If a fair coin is tossed 5 times, what is the probability that we see exactly 3 heads?
a. 0.5000
b. 0.3125
c. 0.8125
d. 0.1875
Answer #1

Probability of head = 0.5

n = 5

P(x=3)

=P ( X =    3   ) = C(5,3) * 0.5^3 * (1-0.5)^2 =            0.3125

b. 0.3125

Please revert back in case of any doubt.

Please upvote. Thanks in advance.